OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (2 points) Part 3: Explain what the solution represents on the graph. (1 point) 3x – 6y = 12 9x + 2y = –24

OpenStudy (anonymous):

Since both equations are of the form y=mx+b, they are both straight lines. You need two points to draw a straight line. 1. So select a value for x and solve for y. That will give you one point to plot on the coordinate plane. 2. Select another value for x and solve for y. That gives you another point to plot on the coordinate plane. 3. Draw a straight line through both points. Repeat steps 1 through 3 for the other equation. Part 2; Rearrange equation one; 3x=12+6y So x=4+2y Substitute into the second equation. 9(4+2y)+2y=-24 So y= -3 Plug y=-3 into the first equation x=4+2(-3) so x=-2 (-2,-3) is the solution Part 3: The point (-2,-3) is the point of intersection between the two lines.

OpenStudy (anonymous):

I suggest you download a free copy of geogebra so you see what it looks like when you graph the two equations.